When carbonyl compounds react with primary amines, an imine is usually produced while a secondary amine gives an enamine. Propose a mechanism to explain this
Answer #1

the reaction mechanisms for the formation of imine and enamine as follows
Ret huis en tuin Ponte - RB Aldehyde CH₃ - Pantinel - 8-c=N-CH, G phone from those . =NCH / Irmine B H - Aldehule & same (R &